# FlippedRAG: Black-Box Opinion Manipulation Adversarial Attacks to Retrieval-Augmented Generation Models

### v3

@@ 5.3. RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
As shown in Table 7, when we manipulated the relevance ranking of documents supporting the target topics, users’ approval of those topics significantly increased. This experiment confirms that opinion manipulation has a substantial effect on users’ perceptions. Comprehensive experimental details of the user study are provided in the Appendix B.4.
@@ 5.4. Mitigation Analyses
We conducted a systematic robustness evaluation of FlippedRAG by implementing multiple defense mechanisms to assess its adversarial resilience. We implemented several defense mechanisms aligned with the core workflow of RAG by selecting phase-specific countermeasures including: detection (filtering) for data, preprocessing defenses for retrieval and generation-phase mitigation.
Given our assumed RAG poisoning scenario involving contamination of publicly indexable web content, we incorporate security detection designed for search engine optimization. Comprehensive technical specifications of the adopted mitigation methods are provided in the Appendix B.4.
@@ 5.4.1. Mitigation Based on Spam Detection
Table 8. TF-IDF based automatic spamicity detection rates (%) of different methods.
@@
Spamicity detection (Zhou and Pei, 2009) is used for discovering term spam. More detail about spam detection is in Appendix B.5.
@@
Under all threshold conditions, the detection rate of triggers generated by FlippedRAG is significantly lower than those generated by PoisonedRAG and exhibits minimal divergence from those observed in clean data. While other baselines demonstrate consistently low detection rates across experimental thresholds, PoisonedRAG exhibited persistently the highest detection rates. This suggests that directly embedding the question in the document lacks sufficient stealth, making it more easily detectable by simple detection methods. The maximal separability between PoisonedRAG’s manipulated documents and clean data occurring at threshold range 0.2 – 0.25. However, achieving a balance between minimizing the risk of erroneously filtering clean data and maintaining high detection efficacy against FlippedRAG remains difficult.
Figure 5. Distributions of log perplexity (PPL) calculated by GPT-2 on clean documents, attacked documents by FlippedRAG, and the baselines.
5.4.2. Mitigation Based on Perplexity Analysis
Perplexity is employed as a metric to evaluate the performance of a probability distribution or probabilistic model in predicting text, assessing whether the text conforms to the conventions of natural language, including coherence and fluency. If a text exhibits a higher perplexity under a language model (LM), it indicates poorer quality. Song et al.(Song et al., 2020) utilized perplexity calculated by GPT-2 in an attempt to distinguish between synthetic text and natural text.
5.4.3. Mitigation Based on Keyword Density
Table 9. Keyword Density(%) of clean documents and documents attacked by FlippedRAG and the baselines.
@@ 5.4.3. Mitigation Based on Keyword Density
Overall Window size
@@
The keyword densitiy of adversarial documents generated by FlippedRAG and other baselines is comparable to that of natural documents in clean data. In contrast, the keyword density of documents generated by PoisonedRAG is significantly higher than that of natural documents, making PoisonedRAG more susceptible to being filtered out by search engine SEO review mechanisms.
@@ 5.4.4. Mitigation Based on Paraphrasing
Cheng et al. (Cheng et al., 2024) and Zou et al. (Zou et al., 2024) attempt to employ paraphrasing defense to test the effectiveness of the RAG attack they proposed.
Given that paraphrasing primarily targets the retrieval phase, the comparative analysis excludes prompt injection attacks and static text manipulation as they exclusively focus on the generation. We also exclude evaluation on GARAG for its inefficacy in opinion manipulation.
The attack effectiveness of FlippedRAG and other baselines against paraphrasing is presented in Table 10. Both the baselines and FlippedRAG exhibit a decline in attack effectiveness when confronted with paraphrasing, yet PoisonedRAG’s and FlippedRAG’s overall attack performance remains notably significant. PoisonedRAG demonstrates a less reduction in attack effectiveness under paraphrasing defense attributing to the direct insertion of queries.
Table 10. Manipulation Effect of the baselines and FlippedRAG against paraphrasing defense. w/o denotes without, w/ denotes with.
@@ 5.4.5. Mitigation Based on Randomized Mask Smoothing
Randomized smoothing is a defense method aiming at achieving certified robustness against adversarial examples. It constructs a smoothed composite function by introducing random variables to the original function (model), thereby enabling the certification of its robustness under certain conditions. We explore randomized mask smoothing as a defense mechanism to counteract RAG attacks in Figure 6.
Figure 6. OMSR(%) under different mask rates and ensemble sizes for FlippedRAG and the baselines.
The experimental results indicate that both PoisonedRAG and FlippedRAG experience a slight decline in attack effectiveness when confronted with randomized masking smoothing. However, since both PoisonedRAG and FlippedRAG employ trigger insertion, the composition of their adversarial documents are similar after masking operations. Additionally, because PoisonedRAG inserts the query itself as the trigger, which enhances retrieval ranking more effectively, the decline in PoisonedRAG’s attack success and opinion manipulation success rate is somewhat smaller.
Our findings further reveal that the success rate of RAG attacks continues to increase when the mask rate ranges from 0% to 30%. This suggests that RAG attacks like FlippedRAG exhibit a certain degree of robustness against random mask smoothing defenses. However, when the mask rate exceeds 50%, the attack success rate shows a more pronounced decline. Nevertheless, an excessively high mask rate may significantly impair the ranking capability of the RAG system, making it difficult for the retrieval model to produce accurate rankings. Consequently, employing random mask smoothing as a defense against FlippedRAG in practical applications is not advisable.
@@ 5.4.6. Mitigation Based on RobustRAG
RobustRAG, proposed by Xiang et al. (Xiang et al., 2024), represents a defense framework specifically targeting retrieval poisoning attacks in RAG systems. It employs an isolate-then-aggregate strategy that leverages the inherent workflow characteristics of RAG architectures to defend against data poisoning attacks.
The manipulation success rate of FlippedRAG against RobustRAG defense is presented in Table 11. Under the RobustRAG mitigation, the OMSR of FlippedRAG decreases compared to scenarios without defensive mechanisms, yet still maintains a success rate reaching approximately 40%—significantly higher than the attack success rates reported by Xiang et al. for factoid questions. This discrepancy arises because RobustRAG was specifically designed as a mitigation method against attacks on factoid and closed-ended questions, where poisoned documents typically contain a single, specific incorrect answer.However, FlippedRAG targets opinion-level manipulation, where a single passage may contain multiple terms or phrases with inherent opinion biases. These opinion-laden terms are quantitatively more prevalent, enabling them to persistently influence the LLM’s output.
@@
Table 11. Manipulation effect of different attacks with RobustRAG defense, where the retrieval model is Contriever.
@@ 5.4.7. Mitigation Analysis Conclusion
FlippedRAG exhibits pronounced adversarial resilience against existing defense mechanisms. In contrast, PoisonedRAG’s reliance on simple heuristic approaches (e.g., keyword saturation) renders it vulnerable to SEO detection methods. Furthermore, its dependence on LLM-generated content introduces measurable deviations from natural text distributions, enabling effective countermeasures through perplexity analysis and RobustRAG mitigation. Other baselines demonstrate limited defensive robustness.
5.5. Parameter Analysis
In the parameter analysis, we investigat the manipulation effectiveness of FlippedRAG when poisoning varying quantities of documents.
Figure 7 demonstrates the opinion manipulation effectiveness of flippedRAG under varying numbers of poisoned documents $N$. As the number of poisoned documents increases from 1 to 10, both the success rate of opinion manipulation and the average stance variation exhibit continuous improvement. When the poisoning quantity $N$ reaches 3 (the default context size $k$ being 3), FlippedRAG already achieves considerable opinion flipping effectiveness. This phenomenon occurs because a larger number of poisoned documents increases the probability of adversarial documents being retrieved into the context. Moreover, when the context size $k$ is 3, three poisoned documents can typically occupy the entire context space in most cases, resulting in a decelerated growth rate of attack effectiveness when the poisoning quantity $N$ exceeds 3.
Figure 7. OMSR(%) and ASV of FlippedRAG with increaing poisoned document number.
@@ 6. Conclusion
In this paper, we proposed FlippedRAG, a transfer-based adversarial attack designed to effectively manipulate the opinion polarity of black-box RAG models, especially on controversial topics. We demonstrated that the underlying retriever of black-box RAG systems can be effectively reverse-engineered via enumeration of critical queries and candidates, which then enables adversaries to construct a precise surrogate retriever. Leveraging this surrogate retriever, we successfully crafted targeted poisoning triggers that effectively manipulated retrieval results and, subsequently, the generated responses. The experimental results demonstrate that our transfer attack strategy effectively alters the opinions expressed in content generated by RAG models. More importantly, opinion manipulation can have profound consequences on users’ cognition, potentially leading them to accept inaccurate or biased information. This could induce cognitive shifts and distort public opinion, especially on controversial topics, highlighting the critical importance of safeguarding against such vulnerabilities in RAG.
@@
This research investigates the security vulnerabilities of RAG systems from the perspective of adversarial opinion manipulation. In accordance with the principles outlined in the Menlo Report—Respect for Persons, Beneficence, Justice, and Respect for Law and Public Interest, as well as the ACM CCS Ethics Guidelines, we undertook a rigorous assessment of the ethical implications associated with our study.
@@
To ensure transparency and reproducibility, all attack implementations, including FlippedRAG, evaluation scripts, and datasets, are hosted on Anonymized GitHub [^3] This paper is under review, so the code has not been open to public yet.. The repository is licensed under the MIT License, allowing reuse while prohibiting malicious applications. We have provided comprehensive documentation, covering environment setup, hyperparameters, and attack pipelines, to ensure independent verification. Sensitive data has been excluded from the repository. All components necessary to replicate our findings are publicly accessible. We are committed to maintaining the repository availability for at least three years and will update the repositories based on community feedback.
@@ A.1. Algorithm
Input: target black-box RAG model $\boldsymbol{\textit{RAG}_{\text{black}}}$, target retrieval model $\boldsymbol{RM}$, surrogate model $\boldsymbol{M_{i}}$, controversial topics $\boldsymbol{Q}$, target topic $q$, target opinion $S_{t}$, corpus $Docs$ , target documents with target opinion $Docs_{t}$, target document $d_{t}$, relevant document $d_{+}$, random sampled document $d_{-}$, poisoned document amount $N$
@@
Please COPY all the given context altogether in [[ ]] including all marks and symbols. Do not omit any sentence of the context”
$i_{2}$ =“Use the following pieces of retrieved context to answer the question…”
// $\textit{RAG}_{\text{black}}$ uses $i_{2}$ as prompt template.
@@
$\boldsymbol{PAT}$: Pairwise Anchor-based Trigger generation strategy.
@@
2 Phase 1. Pairwise Imitation Data Construction and Black-box Retrieval Model Imitation Training
@@ 3       INIT: Dataset $\mathcal{D}\leftarrow\{\}$
4       for $\boldsymbol{q}_{m}\in\boldsymbol{Q}$ do
5             induced rank list $\boldsymbol{R}_{m}{top3}\leftarrow$ $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1};\boldsymbol{Docs})$
             // $\textit{RAG}_{\text{black}}(\boldsymbol{q}_{m}\oplus i_{1})$approx $\boldsymbol{RM(\boldsymbol{q}_{m})}$
6             for $\boldsymbol{d_{+}}_{j}\in\boldsymbol{R}_{m}{top3}$ do
7                   Random sample document from the ranking of origin $\boldsymbol{M_{i}}$ as $\boldsymbol{d_{-}}_{j}$
8                   $\mathcal{D}\leftarrow positive,[\boldsymbol{q}_{m};\boldsymbol{d_{+}}_{j};\boldsymbol{d_{-}}_{j}]$
9                  $\mathcal{D}\leftarrow negative,[\boldsymbol{q}_{m};\boldsymbol{d_{-}}_{j};\boldsymbol{d_{+}}_{j}]$
                   // Reverse $\boldsymbol{d_{+}}_{j}$ and $\boldsymbol{d_{-}}_{j}$ to get the negative triple
10

### v4

@@ 5.3. RQ3: Does opinion manipulation significantly impact users’ perceptions of controversial topics?
In Table 7, opinion polarity of users is quantified by a 7-point Likert scale (1-7), where higher scores indicate stronger support (Pro), while lower scores reflect increasing opposition (Con). For instance, when the target opinion is Con, compared to the control group, greater statistically significant reductions of the opinion polarity of users in the experimental group validate the stronger efficacy of cognitive manipulation. When the target opinion is Pro, compared to the control group, greater increases of the opinion polarity of users in the experimental group validate the stronger efficacy of cognitive manipulation. As shown in Table 7, by manipulating the ranking of documents endorsing specific opinion to bias RAG outputs toward this positions, we induce significant shifts of mean value in the experimental group participants’ opinion polarity. Relative to the control group, these subjects demonstrate cognitive alignment with the target opinion.
Furthermore, we quantified the statistical significance of opinion shift. Formally, we establish the null hypothesis $\texttt{H}_{0}$: There are no significant differences in mean opinion polarity between the experimental and control groups. Subsequent independent sample t-tests conducted on opinions across the three topics yielded p-values below the 0.05 threshold (p < .05 for all comparisons). This necessitates rejection of $\texttt{H}_{0}$, indicating statistically significant differences in opinion polarity between groups and confirming effective cognitive manipulation. This experiment confirms that opinion manipulation has a substantial effect on user perceptions.
@@ 5.4. Mitigation Analyses
We conducted a systematic robustness evaluation of FlippedRAG by implementing multiple defense mechanisms to assess its adversarial resilience. We implemented several defense mechanisms aligned with the core workflow of RAG by selecting phase-specific countermeasures including: detection (filtering) for data, preprocessing defenses for retrieval, and generation-phase mitigation.
Given our assumed RAG poisoning scenario involving contamination of publicly indexable web content, we incorporate security detection designed for search engine optimization.
Figure 5. Distributions of log perplexity (PPL) calculated by GPT-2 on clean documents and attacked documents by FlippedRAG and the baselines.
@@ 5.4.1. Mitigation Based on Spam Detection
Table 8. TF-IDF based automatic spamicity detection rates (%) of different baselines. Lower detection rates correspond to greater attack stealth.
@@
Spamicity detection (zhou2009osd) is used for discovering term spam. More detail about spam detection is in Appendix B.5.
@@
Under all threshold conditions, the detection rate of triggers generated by FlippedRAG is significantly lower than those generated by PoisonedRAG and exhibits minimal divergence from those observed in clean data. While other baselines demonstrate consistently low detection rates across experimental thresholds, PoisonedRAG exhibited persistently the highest detection rates. This suggests that directly embedding the question in the document lacks sufficient stealth, making it more easily detectable by simple detection methods. The maximal separability between the PoisonedRAG’s manipulated documents and clean data occurring at threshold range 0.2 – 0.25. However, achieving a balance between minimizing the risk of erroneously filtering clean data and maintaining high detection efficacy against FlippedRAG remains difficult.
Table 9. Keyword density(%) of clean documents and documents attacked by FlippedRAG and the baselines. High keyword density may result in documents being identified as anomalous and subsequently filtered out.
@@
Overall
Window size
@@
5.4.2. Mitigation Based on Perplexity Analysis
Perplexity is employed as a metric to evaluate the performance of a probability distribution or probabilistic model in predicting text, assessing whether the text conforms to the conventions of natural language, including coherence and fluency. If a text exhibits a higher perplexity under a language model (LM), it indicates poorer quality. Song et al.(song2020adversarial) utilized perplexity calculated by GPT-2 in an attempt to distinguish between synthetic text and natural text.
Figure 6. OMSR(%) under different mask rates and ensemble numbers for FlippedRAG and the baselines.
5.4.3. Mitigation Based on Keyword Density
The keyword density of adversarial documents generated by FlippedRAG and other baselines is comparable to that of natural documents in clean data. In contrast, the keyword density of documents generated by PoisonedRAG is significantly higher than that of natural documents, making PoisonedRAG more susceptible to being filtered out by search engine SEO review mechanisms.
@@ 5.4.4. Mitigation Based on Paraphrasing
Cheng et al. (cheng2024trojanrag) and Zou et al. (zou2024poisonedrag) attempt to employ paraphrasing defense to test the effectiveness of the RAG attack they proposed. The implementation detail of paraphrasing in our experiment is in Appendix B.5.
Given that paraphrasing primarily targets the retrieval phase, the comparative analysis excludes prompt injection attacks and static text manipulation, as they exclusively focus on the generation. We also exclude the evaluation of GARAG for its inefficacy in opinion manipulation.
The attack effectiveness of FlippedRAG and other baselines against paraphrasing is presented in Table 10. Both the baselines and FlippedRAG exhibit a decline in attack effectiveness when confronted with paraphrasing, yet PoisonedRAG’s and FlippedRAG’s overall attack performance remains notably significant. PoisonedRAG demonstrates a less reduction in attack effectiveness under paraphrasing defense attributed to the direct insertion of queries.
Table 10. Manipulation effect of the baselines and FlippedRAG against paraphrasing defense. w/o and w/ denote without and with, respectively.
@@ 5.4.5. Mitigation Based on Randomized Mask Smoothing
Randomized smoothing is a defense method aiming at achieving certified robustness against adversarial examples. It constructs a smoothed composite function by introducing random variables to the original function or the model, enabling the certification of its robustness under certain conditions. We explore randomized mask smoothing as a defense mechanism to counteract RAG attacks in Figure 6. More detail is provided in Appendix B.5.
The experimental results indicate that both PoisonedRAG and FlippedRAG experience a slight decline in attack effectiveness when confronted with randomized masking smoothing. However, since both PoisonedRAG and FlippedRAG employ trigger insertion, the compositions of their adversarial documents are similar after masking operations. Additionally, because PoisonedRAG inserts the query itself as the trigger, which enhances retrieval ranking more effectively, the decline in PoisonedRAG’s attack success and opinion manipulation success rate is somewhat smaller.
Our findings further reveal that the success rate of RAG attacks continues to increase when the mask rate ranges from 0% to 30%. This suggests that RAG attacks like FlippedRAG exhibit a certain degree of robustness against randomized mask smoothing defenses. However, when the mask rate exceeds 50%, the attack success rate shows a more pronounced decline. However, an excessively high mask rate may significantly impair the ranking capability of the RAG system, making it difficult for the retrieval model to produce accurate rankings. Consequently, it is not advisable to employ randomized mask smoothing as a defense against FlippedRAG in practical RAG-like applications.
@@ 5.4.6. Mitigation Based on RobustRAG
RobustRAG, proposed by Xiang et al. (xiang2024certifiably), represents a defense framework specifically targeting retrieval poisoning attacks in RAG systems. It employs an isolate-then-aggregate strategy that leverages the inherent workflow characteristics of RAG architectures to defend against data poisoning attacks. More detail is in Appendix B.5.
The manipulation success rates of FlippedRAG and other baselines against RobustRAG defense is presented in Table 11. Under RobustRAG mitigation, the OMSR of FlippedRAG decreases compared to scenarios without defensive mechanisms, yet still maintains a success rate reaching approximately 40%, significantly higher than the attack success rates reported by Xiang et al. (xiang2024certifiably) for factoid questions. This discrepancy arises because RobustRAG was specifically designed as a mitigation method against attacks on factoid and closed-ended questions, where poisoned documents typically contain a single specific incorrect answer. However, FlippedRAG targets opinion-level manipulation, where a single passage may contain multiple terms or phrases with inherent opinion biases. These opinion-laden terms are quantitatively more prevalent, enabling them to persistently influence the LLM’s output.
@@
Figure 7. OMSR(%) and ASV of FlippedRAG with increasing poisoned document number and larger corpora.
Table 11. Manipulation effect of different attacks against RobustRAG defense, where the retrieval model is Contriever.
@@
5.4.7. Mitigation Based on Privacy Leakage Detection
FlippedRAG employs engineered prompts to induce context disclosure from RAG systems, fundamentally exploiting inherent data leakage vulnerabilities. This approach aligns with prior research by Zeng et al.(zeng2024goodbadexploringprivacy) exploring instruction-based extraction of contextual data from LLMs. To address such leakage vulnerabilities, existing methods mainly leverages LLMs for detection. Two defense methods against such attacks were examined in (naseh2025riddlethisstealthymembership), utilizing GPT-4 and Lakera Guard [^5] https://www.lakera.ai/. Lakera Guard is a commercial AI-native security platform designed to detect malicious prompts.
Our pilot experiments on ChatGPT consistently induced faithful reproduction of retrieved source content, which signifies that many RAG-like applications remain vulnerable to data leakage risks. Furthermore, we use our induction instruction to conduct context extraction experiments against the GPT-4 and Lakera Guard detection, with results detailed in Table 13 at Appendix B.5.6. The analysis reveals that even when these detection methods successfully identify the FlippedRAG malicious instruction previously described, attackers can significantly reduce detection probability by minimally modifying the instructions while preserving their capability to induce LLMs to replicate context data.
Privacy leakage detection based on LLMs and prompt engineering exhibits inherent vulnerabilities, enabling attackers to iteratively refine malicious instructions until identifying blind spots in detection mechanisms or overwriting defensive prompts to achieve context extraction. Consequently, such detection demonstrates limited efficacy against FlippedRAG.
Our research findings provide critical insights for designers of RAG-like applications. Context disclosure vulnerabilities must be prioritized for mitigation. Failure to address these vulnerabilities enables attackers to exploit leaked context data, even in black-box systems, exposing the knowledge of internal components and ultimately enabling manipulation attacks.
@@ 5.4.8. Mitigation Analysis Conclusion
FlippedRAG exhibits pronounced adversarial resilience against existing defense mechanisms. In contrast, PoisonedRAG’s reliance on simple heuristic approaches, e.g., keyword stuffing, renders it vulnerable to SEO detection methods. Furthermore, its dependence on LLM-generated content introduces measurable deviations from natural text distributions, enabling effective countermeasures through perplexity analysis and RobustRAG mitigation. Other baselines demonstrate limited defensive robustness.
5.5. Ablation Studies
In the ablation study, we investigate the manipulation effectiveness of FlippedRAG when poisoning varying quantities of documents and on larger corpora.
The left part in Figure 7 demonstrates the opinion manipulation effectiveness of FlippedRAG with varying numbers of poisoned documents $N$. As the number of poisoned documents increases from 1 to 10, both the success rate of opinion manipulation and the average stance variation exhibit continuous improvement. When the poisoning quantity $N$ reaches 3 (the default context size $k$ being 3), FlippedRAG already achieves considerable opinion flipping effectiveness. This phenomenon occurs because a larger number of poisoned documents increases the probability of adversarial documents being retrieved into the context. Moreover, when the context size $k$ is 3, three poisoned documents can occupy the entire context space in most cases, resulting in a decelerated growth rate of attack effectiveness when the poisoning quantity $N$ exceeds 3.
The right part in Figure 7 presents our attack performance against the RAG based on the embedding model Contriever and the LLM Llama-3 across varying corpus scales. We progressively expand the corpus by incrementally sampling documents from the MS MARCO collection, scaling from 2,000 to 100,000 documents ($2k\rightarrow 100k$). The results reveal an initial declining trend in manipulation efficacy as corpus size increases; however, the magnitude of degradation is small and it diminishes with corpus expansion. Critically, this experiment demonstrates the sustained attack efficacy even at large corpus scales.
@@ 6. Conclusion
In this paper, we proposed FlippedRAG, a transfer-based adversarial attack designed to effectively manipulate the opinion polarity of black-box RAG models, especially on controversial topics. We demonstrated that the underlying retriever of black-box RAG-like systems can be effectively reverse-engineered via enumeration of critical queries and candidates, which then enables adversaries to construct a precise surrogate retriever. Leveraging this surrogate retriever, we successfully crafted targeted poisoning triggers that effectively manipulated the retrieval results and, subsequently, the generated responses. The experimental results demonstrate that our transfer attack strategy effectively alters the opinions expressed in the content generated by the RAG models.
More importantly, opinion manipulation can have profound consequences on users’ cognition, potentially leading them to accept inaccurate or biased information. This could induce cognitive shifts and distort public opinion, especially on controversial topics, highlighting the critical importance of safeguarding against such vulnerabilities in RAG.
@@
This research investigates the security vulnerabilities of Retrieval-Augmented Generation (RAG) systems from the perspective of adversarial opinion manipulation. In accordance with the principles outlined in the Menlo Report—Respect for Persons, Beneficence, Justice, and Respect for Law and Public Interest, as well as the ACM CCS Ethics Guidelines, we undertook a rigorous assessment of the ethical implications associated with our study.
@@
To ensure transparency and reproducibility, all attack implementations, including FlippedRAG, evaluation scripts, and datasets, are hosted on Anonymized GitHub [^6] https://anonymous.4open.science/r/FlippedRAG-BCE5/. The repository is licensed under the MIT License, allowing reuse while prohibiting malicious applications. We have provided comprehensive documentation, covering environment setup, hyperparameters, and attack pipelines, to ensure independent verification. Sensitive data has been excluded from the repository. All components necessary to replicate our findings are publicly accessible. We are committed to maintaining the repository availability for at least three years and will update the repositories based on community feedback.
@@ A.1. Algorithm
Input: target black-box RAG model $\boldsymbol{\textit{RAG}_{\text{black}}}$, target retrieval model $\boldsymbol{RM}$, surrogate model $\boldsymbol{M_{i}}$, controversial topics $\boldsymbol{Q}$, target topic $q$, target opinion $S_{t}$, corpus DocsDocs , target documents with target opinion $Docs_{t}$, target document $d_{t}$, relevant document $d_{+}$ as the positive sample, less relevant document $d_{-}$ as the negative sample, poisoned document amount NN, random sampling amount $\alpha$
@@
Please COPY all the given context altogether…”
@@
$\boldsymbol{PAT}$: Pairwise Anchor-based Trigger generation.
@@
2 Phase 1. Contrastive Data Construction and Black-box Retrieval Model Imitation Training
@@ 3   INIT: Dataset $\mathcal{D}\leftarrow\{\}$
4 for $q_{m}\in\boldsymbol{Q}$ do
5      induced the context with top3 relevant documents $\boldsymbol{RM_{3}(q_{m})}\leftarrow$ $\textit{RAG}_{\text{black}}(q_{m}\oplus i_{1};\boldsymbol{Docs})$
    // $\textit{RAG}_{\text{black}}(q_{m}\oplus i_{1})$ $\approx$ $\boldsymbol{RM_{3}(q_{m})}$
6    for $d_{+_{j}}\in\boldsymbol{RM}_{3}(q_{m})$ do
7       if $j<3$ then
8            sample $d_{-_{m}}$ from lower-ranked positions in $\boldsymbol{RM}_{3}(q_{m})$
9          $\mathcal{D}\leftarrow positive,[q_{m};d_{+_{j}};d_{-_{m}}]$
10          $\mathcal{D}\leftarrow negative,[q_{m};d_{-_{m}};d_{+_{j}}]$